Let X1 and X2 be independent gamma distribution random variables with gamma (a1,1) and gamma (a2, 1). Find the marginal distributions of x1/(x1+x2) and x2/(x1+x2).
Answer #1

Let Y1 = X1/(X1 + X2) and Y2 = X1 + X2

We will use Jacobian method to calculate the joint distributions of x1/(x1+x2) and x1+x2.

the pleat x, and X-ao Taリー20 IS 2- se Can J